OpenStudy (anonymous):

Use the rectangles in each graph to approximate the area of the region bounded by y=sin(x), y=0, x=0, x=pi

OpenStudy (anonymous):

without the graph Im afraid we cannot help

OpenStudy (anonymous):

im trying to get the graphs

OpenStudy (anonymous):

how is that?

OpenStudy (anonymous):

Number one. Lets do this

OpenStudy (anonymous):

So you can see the scale. From 0 to 1 there are 5 equidistant markers. Thus 1/5=0.2 for each scale.

OpenStudy (anonymous):

on the y axis?

OpenStudy (anonymous):

That is on the y axis. On the x axis, first image there is pi/4 scales.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The first triangle goes up to 3.5 scale=3.5*.0.2=0.7 and it goes to the right by pi/4. So the area is pi/4 * 0.7

OpenStudy (anonymous):

What would be the second rectangles area? ( sorry I put triangle before by accident)

OpenStudy (anonymous):

why did you multiply 3.5 by 0.2?

OpenStudy (anonymous):

Because the first rectangles goes up to 3.5 scales. Thus one side is 3.5*0.2 long

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

The other side has length pi/4. The area =0.7 *pi/4 (1st rectangle)

OpenStudy (anonymous):

so the second rectangle is 1*pi/4=pi/4?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

and the first is 0.7pi/4

OpenStudy (anonymous):

And the third is equal to the first. So if you add them up you will get the answer. What is different in the second picture?

OpenStudy (anonymous):

there are 5 rectangles

OpenStudy (anonymous):

Yes, because the scale on the x axis is different. What is it?

OpenStudy (anonymous):

pi/6

OpenStudy (anonymous):

Correct again. I think now you have every tool to finish this question

OpenStudy (anonymous):

also i think the area for the first graph is 3.4pi/4

OpenStudy (anonymous):

im not sure if thats write

OpenStudy (anonymous):

oops right*

OpenStudy (anonymous):

0.7 *pi/4 +1*pi/4+0.7*pi/4=(0.7+1+.0.7)*pi/4=2.4*pi/4

OpenStudy (anonymous):

oh i see where i went wrong

OpenStudy (anonymous):

Probably it is neater to present this solution like this: \[0.6\pi \]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

For your interest the exact solution, so the area under this graph from 0 to pi is equal to 1.

OpenStudy (anonymous):

Oh sorry that was half of it :) It is equal to 2.

OpenStudy (anonymous):

it is?

OpenStudy (anonymous):

We got 0.6pi=1.885 So it is fairly good approximation. If you do the 2nd picture, the result should be closer to 2.

OpenStudy (anonymous):

Yes it is, that area =2

OpenStudy (anonymous):

yes it is :)

OpenStudy (anonymous):

It is a better approximation. If you split the area into 1000000000000 rectangles than the answer might be 1.999999999

OpenStudy (anonymous):

This is where Newton and Leibniz invented Calculus. If you split the graph into infinite many rectangles than the area will be 2!

OpenStudy (anonymous):

well you learn something new everyday

OpenStudy (anonymous):

:)